# Getting Bored of Cyberwar: Exploring the Role of Low-level Cybercrime Actors in the Russia-Ukraine Conflict

### v6

@@ Abstract.
There has been substantial commentary on the role of cyberattacks carried by low-level cybercrime actors in the Russia-Ukraine conflict. We analyse 358k web defacement attacks, 1.7M reflected DDoS attacks, 1 764 Hack Forums posts mentioning the two countries, and 441 announcements (with 58k replies) of a volunteer hacking group for two months before and four months after the invasion. We find the conflict briefly but notably caught the attention of low-level cybercrime actors, with significant increases in online discussion and both types of attack targeting Russia and Ukraine. However, there was little evidence of high-profile actions; the role of these players in the ongoing hybrid warfare is minor, and they should be separated from persistent and motivated ‘hacktivists’ in state-sponsored operations. Their involvement in the conflict appears to have been short-lived and fleeting, with a clear loss of interest in discussing the situation and carrying out both defacement and DDoS attacks against either Russia or Ukraine after a few weeks.
@@ 1. Introduction
Researchers, politicians, and journalists have long been fascinated by ‘cyberwar’ – the spectre of armed conflict between nations spilling over into attacks conducted over the Internet (Rid, 2012). ‘Colder’ forms of inter-state conflict are characterised by espionage and intelligence gathering, which may facilitate the degradation of online systems once hostilities commence (Everett, 2013). Alongside this there has been a thirty-year history of speculation around how the tools and techniques of the cybercrime underground – Distributed Denial of Service (DDoS) attacks, disruption and compromise of services, web defacements, and similar techniques – might allow civilians to play a role in a ‘hot’ war between developed nations (Arquilla and Ronfeldt, 1993). Much of this speculation, drawing from criminological models of low-level cybercrime groups and on links between this underground and well-organised ‘hacktivist’ movements, has argued these groups would play a crucial role, making the future of war hybrid, chaotic, and unpredictable (Wall, 2007). The 2022 invasion of Ukraine provides an opportunity to assess what has happened in practice.
Russia and Ukraine have a long history of electronic information warfare (Margarita Jaitner, 2015) and are among the most active cybercrime hubs (Lusthaus et al., 2020). When Russia invaded Ukraine on 24 February 2022, war-related attacks on the two countries were regularly reported (New York Times, 2022). A popular narrative is that the engagement of low-level cybercrime actors and volunteers could be a game changer and could undermine Russia’s war (Foreign Policy, 2022). Some commentators predicted it will be the first full-scale cyberwar (Atlantic Council, 2022), its effects will last for decades (Serpanos and Komninos, 2022), and youngsters would be drawn into a ‘cyberwar’ by joining IT Army of Ukraine – a group backed by the Ukrainian state to co-ordinate volunteers and civilians to help disrupt Russian assets (Soesanto, 2022; Mykhailo Fedorov, 2022). Some have suggested a real cyber war, predicting hacktivist attacks on Russia would escalate further through 2022 (TASS Russian News Agency, 2022). These narratives regularly appear in the press and play a role in shaping domestic policy responses to cybercrime. Although less likely to grab headlines, a contrary narrative around ‘overhyped cyberwar’ suggests cyber operations in the conflict have been slow (Maschmeyer and Dunn Cavelty, 2022) and insignificant (Kostyuk and Gartzke, 2022), while claims of an unprecedented level of cyberattacks and Russia’s vaunted cyber capabilities are questionable (Nature News Explainer, 2022; Givens et al., 2023; Willett, 2022). GCHQ commented the cyber conflict had not yet materialised (Financial Times, 2022) and pointed to the resilience of Ukraine’s defences (The Economist, 2022).
@@
We longitudinally and statistically measure activities linked with low-level cybercrime actors, including web defacements (§4) and DDoS attacks (§5). The findings are incorporated with analyses of hacking community reactions, including a pro-Ukraine volunteer group (§6). The role of these actors in the conflict is discussed in §7. This study was ethically approved (Appendix §A). All data and scripts are available to researchers on request (Appendix §B).
@@
| Reason and motivation | ||||||
| Type of vulnerability | ||||||
@@ 2. Background and Related Work
Information warfare has long been part of ‘hybrid’ modern conflicts, especially around the control of communications (Hoffman, 2007; Libiseller, 2023). The enemy’s ability to spread news and propaganda can be degraded by targeting crucial sites, public services, broadcast and telecom infrastructure. Censorship is often used during wartime (Price, 1942); governments block access to global services, especially social networks and media platforms, to suppress unwanted narratives. Russia blocked news and anti-war domains when the conflict started (Statista, 2022b; Ramesh et al., 2023), and lost access to foreign sites (Ramesh et al., 2023) and service providers (Jonker et al., 2022). Ukrainian users experienced degraded network performance (Jain et al., 2022), while Ukrainian supporters tried unconventional channels such as online reviews to bypass censorship (Moreno et al., 2023). Attacks are not just online; Russian missiles hit TV towers in Kyiv in early March 2022 (The Guardian, 2022b).
Some links between kinetic warfare and ‘nationalistic’ cyberattacks has been reported. Ukrainian firms were hit by data wipers such as CaddyWiper and NotPetya (ZDNET, 2022; Bleeping Computer, 2022), DDoS attacks (State Sites of Ukraine, 2022; Forbes, 2022b) and phishing campaigns (The Hacker News, 2022); Ukraine supporters have used spam senders to distribute propaganda in Russia (BBC, 2022) and have stolen cryptocurrency from Russian wallets (Forkast News, 2022). Ukrainian universities were hacked (Wordfence, 2022), the Ukrainian electricity grid was hit by Industroyer2 (HeadMind, 2022), and the Ukrainian satellite Internet was downed (Reuters, 2022). Attackers self-identifying with the Anonymous movement declared a ‘cyberwar’ on Russia (The Guardian, 2022a) with attacks against Russian Ministry of Defence databases (Cybernews, 2022b) and state TV channels (Daily Mail, 2022), and Killnet struck back (Metro News, 2022). Russia intermittently received attacks instigated by volunteer hacktivists of the IT Army of Ukraine (Wired, 2022; Foreign Policy, 2022).
@@
One type of attack linked with the low-level cybercrime actors is web defacement (Romagna and van den Hout, 2017), which accounted for around 20% of online attacks in 2014 (Hackmageddon, 2015) and is often organised into discrete campaigns (Maggi et al., 2018). Attackers (or defacers) gain unauthorised access using off-the-shelf tools and simple exploits, then alter sites’ appearance to demonstrate success (Maggi et al., 2018). Defacers have heterogeneous developmental trajectories (van de Weijer et al., 2021); they are often organised in groups (Perkins et al., 2022) and have been using online archives (Kurzmeier, 2020) as a ‘hall of fame’ to show off their achievements to gain reputation. Defacements are mostly hobbies or pranks with greetings to peers (Woo et al., 2004), but some advertise hacking tools or services to make money, or display other motives such as a wish for community recognition, patriotic, religious and political views (Banerjee et al., 2022; Romagna and van den Hout, 2017). Defacement may cause economic harm (Denning, 2011; Andress and Winterfeld, 2011) and has occasionally been used as a proxy for terrorist and other serious activities (Holt et al., 2022). Another simple type of large-scale attack linked to low-level cybercrime actors is amplified DDoS, with many attributed to DDoS-for-hire or ‘booter’ services (Krupp et al., 2017). Such services abound (Karami and McCoy, 2013), and off-the-shelf DDoS tools are widely available; they were tailored and provided to pro-Ukrainian volunteers early in the conflict to attack Russian infrastructure.
@@ 3. Methods and Datasets
We use several quantitative datasets collected regularly and separately, spanning 1 January to 30 June 2022;111 No further substantial changes have been observed beyond the six-month mark. We thus decided to maintain that timeframe, which is sufficient to deliver our narratives. timestamps are normalised to UTC. To determine if the conflict has impacts resulting in different means (or mean ranks) of daily cyberattacks and hacking discussions, we separate the period into three eras; : before the invasion, from 1 January to 24 February 2022; : around one month immediately after the invasion, from 24 February to 31 March 2022; and : from 1 April to 30 June 2022. We then apply unpaired statistical tests, using One-way ANOVA or Kruskal-Wallis depending on the data distribution; the null hypothesis is there is no significant difference between three eras. We use post-hoc tests: Tukey-Kramer for ANOVA or Dunn’s for Kruskal-Wallis to identify pairs causing the changes if any. The effect size is measured by , ranging [0, 1]; : almost no effect; : small effect; : medium effect; : large effect (Miles and Shevlin, 2000).
@@
Web Defacement Attacks. We fully scrape the most popular active defacement archives during the period; see Table 1. We started with Zone-H, the largest and most popular one (since March 2002) providing cybersecurity news and self-reported defacements along with hacking content (Kurzmeier, 2020). We then took out the most active defacers from Zone-H and investigated their online presence, including the archives on which they report attacks. Some defacers promote their attacks on Twitter and Telegram; we also looked there. We then shortlisted the five largest defacement archives by attack volume, including OwnzYou (since January 2021), Zone-Xsec (since May 2020), Haxor-ID (since November 2019), and Defacer-Pro (since June 2021). Smaller archives were historically active (Maggi et al., 2018), but either vanished (Hack Mirror and Mirror Zone) or have hosted different content (Hack-CN and MyDeface). While not all compromised sites get reported, measuring trends from the most reputed archives is likely informative. The country of defaced sites is identified based on ccTLD, IP geolocation, and geolocation of the AS hosting the sites, excluding CDNs (Appendix §C). The defacement submission process is detailed in Appendix §D. We ensure data completeness and bypass challenges e.g., Captcha and IP blocking (Appendix §E).
@@
Further steps are performed to enhance data reliability. First, many on-hold submissions are valid but were never verified; we perform a semi-automatic validation using the messages left on defaced pages (Appendix §F). Second, submissions may be reported to multiple archives to broaden their visibility. We de-duplicate across and within archives by hashing their content (see Appendix §G). Third, as ‘notifier’ can be arbitrary, typos can give a single attacker multiple identities; we correct typos across all archives using handles’ similarity and messages left on defaced pages (see Appendix §G). In total, 137 339 reports were verified by the archives, 97 652 were automatically validated by us and a further 39 972 were validated semi-automatically. 40 799 (11.00%) duplicate reports are merged across all archives. Of the remaining 317 049 reports, we analyse the 274 963 validated submissions (86.73%, around 1 500 per day). Of these, 4 347 defacer handles are also unified to 3 454.
@@
UDP Amplification DDoS Attacks. We use 1.7M DDoS attack records gathered by a honeypot network emulating protocols vulnerable to reflected UDP attacks (Thomas et al., 2017). A flow of packets is considered to be an attack if any sensor observes at least 5 packets for the same victim IP or IP prefix, and the attack is deemed to last from the first packet until the last packet preceding a 15 minute period without further packets. In 2022, the median number of honeypots contributing data was 50, 95% CI [34, 51]; the median number of observed attacks per week was 35 000, 95% CI [11 900, 271 000] and on IP prefixes of 438, 95% CI [0, 3 480]; the median attack duration was 1.53 minutes, while the maximum was 11 300 minutes. The country of victims is identified by geolocation of the IP address and AS hosting it, excluding CDNs (see Appendix §C).
@@
Underground Forum Discussions. Online forums are structured around subforums containing threads with multiple posts. To assess changes in discussion topics within the hacking community, we use a snapshot of the most popular hacking forum, Hack Forums from the CrimeBB dataset (Pastrana et al., 2018b). The forum is a place for users to learn about attacks and trade cybercrime tools and services. Many are low-level actors; however, some have been prosecuted for cybercrime-related activities (Pastrana et al., 2018a). We extract all 123 threads within the 6-month period consisting of at least one post with the keywords ‘Russia’ and/or ‘Ukraine’ (case-insensitive): 115 related to Russia, 108 related to Ukraine, in which 100 related to both. We then use all 1 279 posts from 84 highly relevant threads – those with titles directly having the keywords. For the rest 39 less-relevant threads, we count 485 posts directly consisting of the keywords. In total, 1 764 relevant posts made by 372 users are analysed.
@@
We believe the involved ‘volunteer hacktivists’ are mostly low-level actors, as much of their activity depends on tools provided by the group. We collect 441 announcements with 57 757 replies and 900k emoji reactions posted in the channel from its inception until 30 June 2022 using Telethon, which interacts with official Telegram APIs to fully capture messages and metadata. We then used regular expressions to extract promoted IP addresses and domains; subdomains such as www.xyz.ru and smtp.xyz.ru are combined. Besides Russian and Belarusian domains (.ru, .su, .by), top level domains (e.g., .tv, .com) are also targeted. URL shorteners (e.g., goo.gl) and online services (e.g., youtube.com) are excluded, resulting in 3 845 targets: 2 291 IP addresses (59.58%) and 1 554 domains (40.42%).

### v7

@@ Abstract.
There has been substantial commentary on the role of cyberattacks carried out by low-level cybercrime actors in the Russia-Ukraine conflict. We analyse 358k website defacement attacks, 1.7M UDP amplification DDoS attacks, 1 764 posts made by 372 users on Hack Forums mentioning the two countries, and 441 Telegram announcements (with 58k replies) of a volunteer hacking group for two months before and four months after the invasion. We find the conflict briefly but notably caught the attention of low-level cybercrime actors, with significant increases in online discussion and both types of attacks targeting Russia and Ukraine. However, there was little evidence of high-profile actions; the role of these players in the ongoing hybrid warfare is minor, and they should be separated from persistent and motivated ‘hacktivists’ in state-sponsored operations. Their involvement in the conflict appears to have been short-lived and fleeting, with a clear loss of interest in discussing the situation and carrying out both website defacement and DDoS attacks against either Russia or Ukraine after just a few weeks.
@@ 1. Introduction
Researchers, politicians, and journalists have long been fascinated by ‘cyberwar’ – the spectre of armed conflict between nations spilling over into attacks conducted over the Internet (Rid, 2012). ‘Colder’ forms of inter-state conflict are characterised by espionage and intelligence gathering, which may facilitate the degradation of online systems once hostilities commence (Everett, 2013). Alongside this, there has been a thirty-year history of speculation around how the tools and techniques of the cybercrime underground – Distributed Denial of Service (DDoS) attacks, disruption and compromise of services, web defacements, and similar techniques – might allow civilians to play a role in a ‘hot’ war between developed nations (Arquilla and Ronfeldt, 1993). Much of this speculation, drawing from criminological models of low-level cybercrime groups and on links between this underground and well-organised ‘hacktivist’ movements, has argued these groups would play a crucial role, making the future of war hybrid, chaotic, and unpredictable (Wall, 2007). The 2022 Russian invasion of Ukraine provides an opportunity to assess what has happened in practice.
Russia and Ukraine have a long history of electronic information warfare (Margarita Jaitner, 2015) and are among the most active cybercrime hubs (Lusthaus et al., 2020). When Russia invaded Ukraine on 24 February 2022, war-related attacks on the two countries were regularly reported (New York Times, 2022). A popular narrative is that the engagement of low-level cybercrime actors and volunteers could be a game changer and could undermine Russia’s war (Foreign Policy, 2022). Some commentators predicted it will be the first full-scale cyberwar (Atlantic Council, 2022), its effects will last for decades (Serpanos and Komninos, 2022), and youngsters would be drawn into a ‘cyberwar’ by joining IT Army of Ukraine – a group backed by the Ukrainian state to co-ordinate volunteers and civilians to help disrupt Russian assets (Soesanto, 2022; Mykhailo Fedorov, 2022). Some have suggested a ‘real cyberwar’, predicting hacktivist attacks on Russia would escalate further through 2022 (TASS Russian News Agency, 2022). These narratives regularly appear in the press and play a role in shaping domestic policy responses to cybercrime. Although less likely to grab headlines, a contrary narrative around ‘overhyped cyberwar’ suggests cyber operations in the conflict have been slow (Maschmeyer and Dunn Cavelty, 2022) and insignificant (Kostyuk and Gartzke, 2022), while claims of an unprecedented level of cyberattacks and Russia’s much-vaunted cyber capabilities are questionable (Nature News Explainer, 2022; Givens et al., 2023; Willett, 2022). GCHQ commented the cyber conflict had not yet materialised (Financial Times, 2022) and pointed to the resilience of Ukraine’s defences (The Economist, 2022).
@@
We longitudinally and statistically measure activities linked with low-level cybercrime actors, including web defacements (§4) and DDoS attacks (§5). The findings are incorporated with analyses of hacking community reactions, including a pro-Ukraine volunteer group (§6). The role of these actors in the conflict is discussed in §7. This study was ethically approved (see Appendix §A). All data and scripts are available to researchers on request (see Appendix §B).
@@
| Reasons and/or motivations | ||||||
| Types of vulnerability | ||||||
@@ 2. Background and Related Work
Information warfare has long been part of ‘hybrid’ modern conflicts, especially around the control of communications (Hoffman, 2007; Libiseller, 2023). The enemy’s ability to spread news and propaganda can be degraded by targeting crucial sites, public services, broadcast and telecom infrastructure. Censorship is often used during wartime (Price, 1942); governments block access to global services, especially social networks and media platforms, to suppress unwanted narratives. Russia blocked news and anti-war domains when the conflict started (Statista, 2022b; Ramesh et al., 2023), and lost access to foreign sites (Ramesh et al., 2023) as well as service providers (Jonker et al., 2022). Ukrainian users experienced degraded network performance (Jain et al., 2022), while Ukrainian supporters tried unconventional channels such as online reviews to bypass censorship (Moreno et al., 2023). Attacks are not just online; Russian missiles hit TV towers in Kyiv in early March 2022 (The Guardian, 2022b).
Some associations between kinetic warfare and ‘nationalistic’ cyberattacks have been reported. Ukrainian firms were hit by data wipers such as CaddyWiper and NotPetya (ZDNET, 2022; Bleeping Computer, 2022), DDoS attacks (State Sites of Ukraine, 2022; Forbes, 2022b) and phishing campaigns (The Hacker News, 2022); Ukraine supporters have used spam senders to distribute propaganda in Russia (BBC, 2022) and have stolen cryptocurrency from Russian wallets (Forkast News, 2022). Ukrainian universities were hacked (Wordfence, 2022), the Ukrainian electricity grid was hit by Industroyer2 (HeadMind, 2022), and the Ukrainian satellite Internet was downed (Reuters, 2022). Attackers self-identifying with the Anonymous movement declared a ‘cyberwar’ on Russia (The Guardian, 2022a) with attacks against Russian Ministry of Defence databases (Cybernews, 2022b) and state TV channels (Daily Mail, 2022), while Killnet struck back (Metro News, 2022). Russia intermittently received attacks instigated by volunteer hacktivists of the IT Army of Ukraine (Wired, 2022; Foreign Policy, 2022).
@@
One type of attack linked with the low-level cybercrime actors is website defacement (Romagna and van den Hout, 2017), which accounted for around 20% of online attacks in 2014 (Hackmageddon, 2015) and is often organised into discrete campaigns (Maggi et al., 2018). Attackers (or defacers) gain unauthorised access using off-the-shelf tools and simple exploits, then alter sites’ appearance to demonstrate success (Maggi et al., 2018). Defacers have heterogeneous developmental trajectories (van de Weijer et al., 2021); they are often organised in groups (Perkins et al., 2022) and have been using online archives (Kurzmeier, 2020) as a ‘hall of fame’ to show off their achievements to gain reputation. Defacements are mostly hobbies or pranks with greetings to peers (Woo et al., 2004), but some advertise hacking tools or services to make money or display other motives such as a wish for community recognition, patriotic, religious and political views (Banerjee et al., 2022; Romagna and van den Hout, 2017). Web defacement may cause economic harm (Denning, 2011; Andress and Winterfeld, 2011) and has occasionally been used as a proxy for terrorist and other serious activities (Holt et al., 2022). Another simple type of large-scale attack linked to low-level cybercrime actors is amplified DDoS, with many attributed to DDoS-for-hire or ‘booter’ services (Krupp et al., 2017). Such services abound (Karami and McCoy, 2013), and off-the-shelf DDoS tools are widely available; they were tailored and provided to pro-Ukrainian volunteers early in the conflict to attack Russian infrastructure and assets.
@@ 3. Methods and Datasets
We use several quantitative datasets collected regularly and separately, spanning 1 January to 30 June 2022;111 No further substantial changes have been observed beyond the six-month mark. We thus decided to maintain that timeframe, which is sufficient to deliver our narratives. timestamps are normalised to UTC. To determine if the conflict has impacts resulting in different means (or mean ranks) of daily cyberattacks and hacking discussions, we separate the period into three eras; : before the invasion, from 1 January to 24 February 2022; : around one month immediately after the invasion, from 24 February to 31 March 2022; and : from 1 April to 30 June 2022. We then apply unpaired statistical tests, using One-way ANOVA or Kruskal-Wallis depending on the data distribution; the null hypothesis is there is no significant difference between the three eras. We use post-hoc tests: Tukey-Kramer for ANOVA or Dunn’s for Kruskal-Wallis to identify pairs causing changes if any. Effect sizes are measured by , ranging [0, 1]; : no effect; : small effect; : medium effect; : large effect (Miles and Shevlin, 2000).
@@
Web Defacement Attacks. We fully scrape the most popular active defacement archives during the period; see Table 1. We started with Zone-H, the largest and most popular one (since March 2002) providing cybersecurity news and self-reported defacements along with hacking content (Kurzmeier, 2020). We then took out the most active defacers from Zone-H and investigated their online presence, including the archives on which they report attacks. Some defacers promote their attacks on Twitter and Telegram; we also looked there. We then shortlisted the five largest defacement archives by attack volume, including OwnzYou (since January 2021), Zone-Xsec (since May 2020), Haxor-ID (since November 2019), and Defacer-Pro (since June 2021). Smaller archives were historically active (Maggi et al., 2018), but either vanished (Hack Mirror and Mirror Zone) or have hosted different content (Hack-CN and MyDeface). While not all compromised sites get reported, measuring trends from the most reputed archives is likely indicative. The country of defaced sites is identified based on ccTLD, IP geolocation, and geolocation of the AS hosting the sites, excluding CDNs (Appendix §C). The defacement submission process is detailed in Appendix §D. We ensure data completeness and bypass challenges e.g., Captcha and IP blocking (Appendix §E).
@@
Further steps are performed to enhance data reliability. First, many on-hold submissions are valid but were never verified; we perform a semi-automatic validation using the messages left on defaced pages (see Appendix §F). Second, submissions may be reported to multiple archives to broaden their visibility. We de-duplicate across and within archives by hashing their content (see Appendix §G). Third, as ‘notifier’ can be arbitrary, typos can give a single attacker multiple identities; we correct typos across all archives using handles’ similarity and messages left on defaced pages (see Appendix §G). In total, 137 339 reports were verified by the archives, 97 652 were automatically validated by us and a further 39 972 were validated semi-automatically; 40 799 (11.00%) duplicate reports are merged across all archives. Of the remaining 317 049 reports, we analyse 274 963 validated submissions (86.73%, around 1 500 per day). Of these, 4 347 defacer handles are also unified to 3 454.
@@
UDP Amplification DDoS Attacks. We use 1.7M DDoS attack records gathered by a honeypot network emulating protocols vulnerable to reflected UDP attacks (Thomas et al., 2017). A flow of packets is considered to be an attack if any sensor observes at least five packets for the same victim IP or IP prefix, and the attack is deemed to last from the first packet until the last packet preceding 15 minutes without further packets. In 2022, the median number of honeypots contributing data was 50, 95% CI [34, 51]; the median number of observed attacks per week was 35 000, 95% CI [11 900, 271 000] and on IP prefixes of 438, 95% CI [0, 3 480]; the median attack duration was 1.53 minutes, while the maximum was 11 300 minutes. The country of victims is identified by geolocation of the IP address and AS hosting it, excluding CDNs (see Appendix §C).
@@
Underground Forum Discussions. Online forums are structured around subforums containing threads with multiple posts. To assess changes in discussion topics within the hacking community, we use a snapshot of the most popular hacking forum, Hack Forums from the CrimeBB dataset (Pastrana et al., 2018b). The forum is a place for users to learn about attacks and trade cybercrime tools and services. Many are low-level actors; however, some have been prosecuted for cybercrime-related activities (Pastrana et al., 2018a). We extract all 123 threads within the six months consisting of at least one post with the keywords ‘Russia’ and/or ‘Ukraine’ (case-insensitive): 115 related to Russia, 108 related to Ukraine, in which 100 related to both. We then use all 1 279 posts from 84 highly relevant threads – those with titles directly having the keywords. For the rest 39 less-relevant threads, we count 485 posts directly consisting of the keywords. In total, 1 764 relevant posts made by 372 active forum users are analysed.
@@
We believe the involved ‘volunteer hacktivists’ are mostly low-level actors, as much of their activity depends on tools provided by the group. We collect 441 announcements with 57 757 replies and 900k emoji reactions posted in the channel from its inception until 30 June 2022 using Telethon, which interacts with official Telegram APIs to fully capture messages and metadata. We then used regular expressions to extract promoted IP addresses and domains; subdomains such as www.xyz.ru and smtp.xyz.ru are combined. Besides Russian and Belarusian domains (.ru, .su, .by), top-level domains (e.g., .tv, .com) are also targeted. URL shorteners (e.g., goo.gl) and online services (e.g., youtube.com) are excluded, resulting in 3 845 targets: 2 291 IP addresses (59.58%) and 1 554 domains (40.42%).